$SPAD/src/input function.input

The Axiom Team December 3, 2016

Abstract

Contents

2 Input for page RationatFunctionPage

— * —

)set break resume
)spool function.output
)set message test on
)set message auto off
)clear all
--S 1 of 33
f := (x - y) / (x + y)
--R
--R        - y + x
--R   (1)  -------
--R         y + x
--R                                          Type: Fraction(Polynomial(Integer))
--E 1
--S 2 of 33
numer f
--R
--R   (2)  - y + x
--R                                                    Type: Polynomial(Integer)
--E 2
--S 3 of 33
denom f
--R
--R   (3)  y + x
--R                                                    Type: Polynomial(Integer)
--E 3
--S 4 of 33
eval(f, x = 1/x)
--R
--R        - x y + 1
--R   (4)  ---------
--R         x y + 1
--R                                          Type: Fraction(Polynomial(Integer))
--E 4
--S 5 of 33
eval(f, [x = y, y = x])
--R
--R        y - x
--R   (5)  -----
--R        y + x
--R                                          Type: Fraction(Polynomial(Integer))
--E 5

———-

Input for page AlgebraicFunctionPage — * —

)clear all
--S 6 of 33
f := sqrt(1 + x ** (1/3))
--R
--R         +--------+
--R         |3+-+
--R   (1)  \|\|x  + 1
--R                                                    Type: Expression(Integer)
--E 6
--S 7 of 33
y := rootOf(y**3 + y**2 - x*y + x**3 - 1, y)
--R
--R   (2)  y
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 33
differentiate(y, x)
--R
--R                 2
--R           y - 3x
--R   (3)  ------------
--R          2
--R        3y  + 2y - x
--R                                                    Type: Expression(Integer)
--E 8
--S 9 of 33
(y + 1) ** 3
--R
--R          2               3
--R   (4)  2y  + (x + 3)y - x  + 2
--R                                                    Type: Expression(Integer)
--E 9
--S 10 of 33
g := inv f
--R
--R             1
--R   (5)  -----------
--R         +--------+
--R         |3+-+
--R        \|\|x  + 1
--R                                                    Type: Expression(Integer)

--E 10
--S 11 of 33
ratPoly g
--R
--R                6     4     2
--R   (6)  (x + 1)?  - 3?  + 3?  - 1
--R                        Type: SparseUnivariatePolynomial(Expression(Integer))
--E 11

———-

Input for page ElementaryFunctionPage — * —

)clear all
--S 12 of 33
f := x * log y * sin(1/(x+y))
--R
--R                      1
--R   (1)  x log(y)sin(-----)
--R                    y + x
--R                                                    Type: Expression(Integer)
--E 12
--S 13 of 33
eval(f, [x = y, y = x])
--R
--R                      1
--R   (2)  y log(x)sin(-----)
--R                    y + x
--R                                                    Type: Expression(Integer)
--E 13
--S 14 of 33
eval(f, log y = acosh(x + sqrt y))
--R
--R                1          +-+
--R   (3)  x sin(-----)acosh(\|y  + x)
--R              y + x
--R                                                    Type: Expression(Integer)
--E 14

———-

Input for page FunctionSimplificationPage — * —

)clear all

--S 15 of 33
f := cos(x)/sec(x) * log(sin(x)**2/(cos(x)**2+sin(x)**2))
--R
--R                             2
--R                       sin(x)
--R        cos(x)log(-----------------)
--R                        2         2
--R                  sin(x)  + cos(x)
--R   (1)  ----------------------------
--R                   sec(x)
--R                                                    Type: Expression(Integer)
--E 15
--S 16 of 33
g := simplify f
--R
--R              2            2
--R   (2)  cos(x) log(- cos(x)  + 1)
--R                                                    Type: Expression(Integer)
--E 16
--S 17 of 33
h := sin2csc cos2sec g
--R
--R                  2
--R            sec(x)  - 1
--R        log(-----------)
--R                    2
--R              sec(x)
--R   (3)  ----------------
--R                   2
--R             sec(x)
--R                                                    Type: Expression(Integer)
--E 17
--S 18 of 33
expandLog h
--R
--R                  2
--R        log(sec(x)  - 1) - 2log(sec(x))
--R   (4)  -------------------------------
--R                          2
--R                    sec(x)
--R                                                    Type: Expression(Integer)
--E 18
--S 19 of 33
f1 := sqrt((x+1)**3)
--R
--R         +-----------------+

--R         | 3     2
--R   (5)  \|x  + 3x  + 3x + 1
--R                                                    Type: Expression(Integer)
--E 19
--S 20 of 33
rootSimp f1
--R
--R                +-----+
--R   (6)  (x + 1)\|x + 1
--R                                                    Type: Expression(Integer)
--E 20
--S 21 of 33
g1 := sin(x + cos x)
--R
--R   (7)  sin(cos(x) + x)
--R                                                    Type: Expression(Integer)
--E 21
--S 22 of 33
g2 := complexElementary g1
--R
--R                              +---+ 2               +---+          2
--R                    +---+   x\|- 1         +---+  x\|- 1     +---+
--R                   \|- 1 (%e       )  + 2x\|- 1 %e        + \|- 1
--R                   -----------------------------------------------
--R                                           +---+
--R                                         x\|- 1
--R           +---+                      2%e                               +---+
--R        - \|- 1 (%e                                               )  + \|- 1
--R   (8)  ---------------------------------------------------------------------
--R                                +---+ 2               +---+
--R                      +---+   x\|- 1         +---+  x\|- 1     +---+
--R                     \|- 1 (%e       )  + 2x\|- 1 %e        + \|- 1
--R                     -----------------------------------------------
--R                                             +---+
--R                                           x\|- 1
--R                                        2%e
--R                  2%e
--R                                                    Type: Expression(Integer)
--E 22
--S 23 of 33
trigs g2
--R
--R   (9)  sin(cos(x) + x)
--R                                                    Type: Expression(Integer)
--E 23

--S 24 of 33
h1 := sinh(x + cosh x)
--R
--R   (10)  sinh(cosh(x) + x)
--R                                                    Type: Expression(Integer)
--E 24
--S 25 of 33
h2 := realElementary h1
--R
--R               x 2        x     2
--R            (%e )  + 2x %e  + 1
--R            -------------------
--R                       x
--R                    2%e
--R         (%e                   )  - 1
--R   (11)  ----------------------------
--R                  x 2        x
--R               (%e )  + 2x %e  + 1
--R               -------------------
--R                          x
--R                       2%e
--R            2%e
--R                                                    Type: Expression(Integer)
--E 25
--S 26 of 33
htrigs h2
--R
--R   (12)  sinh(cosh(x) + x)
--R                                                    Type: Expression(Integer)
--E 26

———-

Input for page PatternMatchingPage — * —

)clear all
--S 27 of 33
groupSqrt := _rule(sqrt(a) * sqrt(b), sqrt(a*b))
--R
--R           +-+ +-+       +---+
--I   (1)  %B\|a \|b  == %B\|a b
--R                       Type: RewriteRule(Integer,Integer,Expression(Integer))
--E 27
--S 28 of 33
a := sqrt(2) * sqrt(3)

--R
--R         +-+ +-+
--R   (2)  \|2 \|3
--R                                                        Type: AlgebraicNumber
--E 28
--S 29 of 33
groupSqrt a
--R
--R         +-+
--R   (3)  \|6
--R                                                    Type: Expression(Integer)
--E 29
--S 30 of 33
a := (sqrt(x) + sqrt(y))**4
--R
--R                  +-+ +-+    2           2
--R   (4)  (4y + 4x)\|x \|y  + y  + 6x y + x
--R                                                    Type: Expression(Integer)
--E 30
--S 31 of 33
groupSqrt a
--R
--R                  +---+    2           2
--R   (5)  (4y + 4x)\|x y  + y  + 6x y + x
--R                                                    Type: Expression(Integer)
--E 31
--S 32 of 33
sinCosExpand := rule
  sin(-x)    == - sin(x)
  cos(-x)    == cos(x)
  sin(x + y) == sin(x) * cos(y) + sin(y) * cos(x)
  cos(x + y) == cos(x) * cos(y) - sin(x) * sin(y)
  sin((n | integer? n and n > 1) * x) ==_
       sin(x) * cos((n-1)*x) + sin((n-1)*x) * cos(x)
  cos((n | integer? n and n > 1) * x) ==_
       cos(x) * cos((n-1)*x) - sin(x) * sin((n-1)*x)
--R
--R
--R   (6)
--I   {- %BC sin(x) == - %BC sin(x), cos(x) == cos(x),
--R    sin(y + x) == cos(x)sin(y) + cos(y)sin(x),
--R    cos(y + x) == - sin(x)sin(y) + cos(x)cos(y),
--R    sin(n x) == cos(x)sin((n - 1)x) + cos((n - 1)x)sin(x),
--R    cos(n x) == - sin(x)sin((n - 1)x) + cos(x)cos((n - 1)x)}
--R                           Type: Ruleset(Integer,Integer,Expression(Integer))
--E 32

--S 33 of 33
sinCosExpand(sin(x+y-2*z) * cos y)
--R
--R   (7)  - cos(y)sin(2z - y - x)
--R                                                    Type: Expression(Integer)
--E 33
)spool
)lisp (bye)

———-

References^{[1] nothing}

11